OpenStudy (emyoung0):

True or False? The two functions f and g defined by f(x) = 3x + 3 for x real and g(t) = 3t + 3 for t real and positive are equal?

OpenStudy (xguardians):

I believe it's true, you are just changing the variable name?

OpenStudy (emyoung0):

Well, two functions can only be equal if their rules are equal and their domains are the same, which isnt true in this case.. Im leaning towards false. But I guess I'll go with true.

zepdrix (zepdrix):

Well the rule is certainly the same in both cases, multiply by 3, and then add 3. The domains appear to be equal as well, yes? :) `real x`, `real t`

OpenStudy (mazmaz1):

they are essentially the same but strictly speaking about the domain itself they are not rather than the values it wont but all the values do mach up as identical

OpenStudy (mazmaz1):

real but not fully identical although values add up

zepdrix (zepdrix):

Hmm I don't understand what you're saying ^

zepdrix (zepdrix):

Oh I misread part of it. Is the `and positive` only for the t?

OpenStudy (emyoung0):

Must be...

zepdrix (zepdrix):

ugh, poorly worded :c that's unfortunate... If x is all real numbers, then \(\large\rm x\in\left(-\infty,\infty\right)\) If t is all real numbers and positive then \(\large\rm t\in\left[0,\infty\right)\) Different domains, ya? Sorry that wording tricked me :c

OpenStudy (emyoung0):

You're not the only one! Thanks!